OpenStudy (australopithecus):

Do I just use telescoping method to find the sum of the series \[\sum_{n=1}^{\infty}-3/n(n+4)\]

OpenStudy (kinggeorge):

Is that \[\sum_{n=1}^\infty \frac{-3}{n(n+4)}?\]

OpenStudy (australopithecus):

yes how do you do that in latex btw

OpenStudy (kinggeorge):

\sum_{n=1}^\infty \frac{-3}{n(n+4)}

OpenStudy (australopithecus):

thanks, would I use telescoping method, where I use partial fractions then cancel terms and then take the limit

OpenStudy (kinggeorge):

That seems like it would work well.

OpenStudy (australopithecus):

By taking the limit do I receive the sum of the function?

OpenStudy (kinggeorge):

No. If you take the limit, you get 0, but the sum is not 0.

OpenStudy (australopithecus):

right but if I take the limit of the terms that cancel out after using partial fractions

OpenStudy (kinggeorge):

After using partial fractions, you'll get some terms in the beginning that don't cancel. Add these together. Then take the limit of partial sum expansion. Add the limit, and any other terms at the end that don't entirely cancel.

OpenStudy (australopithecus):

right and that will give me the sum right :)

OpenStudy (kinggeorge):

Right. Feel free to post any work here and I can check it.

OpenStudy (australopithecus):

ok thanks

OpenStudy (australopithecus):

also would you mind checking this I determine the series was convergent based on the comparison test  where the series 1/r^(p) is convergent when p>1 therefore by C.T. the original series is converget